At the time she was hired as a server at the Grumney Family Restaurant, Beth Brigden was told, “You can average $83 a day in tips.” Assume the population of daily tips is normally distributed with a standard deviation of $4.07. Over the first 45 days she was employed at the restaurant, the mean daily amount of her tips was $84.86. At the 0.05 significance level, can Ms. Brigden conclude that her daily tips average more than $83?

Compute the value of the test statistic. (Round your answer to 2 decimal places.)

What is the p-value? (Round your answer to 4 decimal places.)
Answer #1

Solution :

The null and alternative hypothesis is ,

H0 :   = 83

Ha : > 83

Test statistic = z =

= ( - ) / / n

= (84.46 - 83) / 4.07 / 45

Test statistic = z = 2.41

This is the right tailed test,

P(Z > 2.41 ) = 1-P (Z < 2.41) = 1 - 0.9920

P-value = 0.0080